# Group Therapy for Halos: Advancing Halo Mass Estimation for Galaxy Groups

### v1

@@ 2.1.1 SHARK
Shark v2.0 (Lagos et al., 2018, 2024) is an open-source, modular semi-analytic model of galaxy formation and evolution. The latest version incorporates significant advancements, including an improved treatment of angular momentum evolution, several environmental processes including ram pressure and tidal stripping, and updated feedback models. The free parameters in Shark are calibrated to reproduce the observed stellar mass function, star formation rate density, and cold gas scaling relations at $z=0$.
@@
The Shark runs analysed in this work are based on the SURFS suite of $N$-body simulations (Elahi et al., 2018), specifically medi-SURFS, which spans $(210~h^{-1}~\mathrm{Mpc})^{3}$ with $1536^{3}$ dark matter particles, yielding a particle mass of $2.21\times 10^{8}h^{-1}M_{\odot}$. Haloes, sub-haloes, and merger trees are constructed using HBT+HERONS (Chandro-Gómez et al., 2025). The Shark simulated light cones used in this work correspond to WAVES WIDE (North + South) light cones, totalling $\sim 1,100\,\mathrm{deg}^{2}$ in area. These light cones were constructed using the pipeline described in Lagos et al. (2019): the survey geometry and magnitude selections are built using Stingray (Chauhan et al., 2019), and the galaxy SEDs are built using ProSpect (Robotham et al., 2020). A lower stellar mass limit of $\log M_{\star}>7.5~[M_{\odot}]$ was applied to the mock light cones to ensure completeness and reliability in the resulting galaxy sample. The simulation adopts a Planck 2015 cosmology (Planck Collaboration et al., 2016), with $\Omega_{\mathrm{m}}=0.3121$, $\Omega_{\Lambda}=0.6879$, $\Omega_{\mathrm{b}}=0.0491$, $h=0.6751$, $\sigma_{8}=0.8150$, and $n_{\mathrm{s}}=0.9653$.
2.1.2 SAGE
@@
The Semi-Analytic Galaxy Evolution (SAGE) model (Croton et al., 2016) is a flexible, publicly available semi-analytic framework, building upon the Munich model lineage (Croton et al., 2006). SAGE incorporates detailed prescriptions for radiative cooling, star formation, stellar and AGN feedback, black hole growth, and environmental processes. The model parameters are tuned to match the observed stellar mass function and galaxy colour distributions at $z=0$.
@@
For this study, SAGE is applied to the BOLSHOI $N$-body simulation (Klypin et al., 2011), which covers a volume of $(250~h^{-1}~\mathrm{Mpc})^{3}$ with $2048^{3}$ particles, corresponding to a mass resolution of $1.35\times 10^{8}~h^{-1}~M_{\odot}$. The SAGE simulated light cones used in this work consisted of 10 lightcones, each with an area of $\sim 1,960\,\mathrm{deg}^{2}$. Haloes are identified using the ROCKSTAR phase-space halo finder (Behroozi et al., 2013a), and merger trees are constructed with the Consistent Trees algorithm (Behroozi et al., 2013b). A lower stellar mass limit of $\log M_{\star}>7.5~[M_{\odot}]$ was applied to the mock light cones to ensure completeness and reliability in the resulting galaxy sample. The adopted cosmology is based on WMAP7 (Komatsu et al., 2011), with $\Omega_{\mathrm{m}}=0.270$, $\Omega_{\Lambda}=0.730$, $\Omega_{\mathrm{b}}=0.0469$, $h=0.70$, $\sigma_{8}=0.82$, and $n_{\mathrm{s}}=0.95$.
2.1.3 GAEA
@@
The Galaxy Evolution and Assembly (GAEA) semi-analytic model (De Lucia et al., 2014; Hirschmann et al., 2016; De Lucia et al., 2024) provides an independent theoretical benchmark in this work. The latest version includes updated prescriptions for AGN feedback, environmental processes affecting satellites, black hole accretion, disk instabilities, and starburst activity (Fontanot et al., 2020; De Lucia et al., 2011).
@@
2.2 SGP: Southern Galactic Pole Catalogue
The Southern Galactic Pole (SGP) observational dataset, introduced in Van Kempen et al. (2024), provides a comprehensive spectroscopic and photometric dataset for galaxies at $z<0.1$ over 376 deg2 ($340^{\circ}<\mathrm{RA}<26^{\circ}$, $35.3^{\circ}<\mathrm{Dec}<-25.8^{\circ}$), overlapping the 2-degree Field Galaxy Redshift Survey (2dFGRS) and GAMA G23 fields (Colless et al., 2001; Driver et al., 2009). The sample combines redshifts primarily from 2dFGRS and GAMA, supplemented by 2-degree Field Lensing Survey (2dFLenS), the 6-degree Field Galaxy Redshift Survey (6dFGRS), the 2MASS Redshift Survey (2MRS), and the Million Quasars catalogue (MILLIQUAS; Blake et al., 2016; Jones et al., 2004; Macri et al., 2019; Flesch, 2021), yielding 24,656 unique spectroscopic sources. This results in a heterogeneous dataset, with 22,933 ($\sim$93%) sources cross-matched to photometry from the Wide-Field Infrared Survey Explorer (WISE; Wright et al., 2010).
Given the heterogeneity of the parent sample, homogeneous sub-samples are defined for robust statistical analyses of the stellar mass function (SMF) and halo mass function (HMF). The SGP–2dF sample comprises galaxies with 2dFGRS redshifts across the entire SGP region, ensuring uniform selection. Within the GAMA G23 field, two further sub-samples are constructed: G23–2dF, which includes only 2dFGRS measured galaxies, and G23–GAMA, which utilises the more complete GAMA spectroscopic data. The G23–GAMA sample achieves approximately 2.3 times higher completeness than the G23–2dF sub-sample, providing a valuable benchmark for completeness and systematic effects.
Galaxy groups were identified for the SGP dataset via a graph-theory derived Friends-of-Friends (FoF) algorithm (pyFoF; Lambert et al., 2020)111pyFoF can be acsessed via: https://github.com/TrystanScottLambert/pyFoF, see Van Kempen et al. (2024) for full details on the optimisation of the FoF based finder. The final group catalogue included 1,413 unique groups containing 8,980 galaxies, of which 8,454 were successfully cross-matched with WISE. Cross-matched WISE photometry allows for accurate measurements of stellar masses, estimated via the W1-based relations from Jarrett et al. (2023), and star formation rates, derived from W3 and W4 luminosities following the calibrations of Cluver et al. (2025).
@@ 3 Halo Mass Relations
In the group and cluster regime, halo mass estimates are traditionally derived from dynamical tracers, such as the velocity dispersion of member galaxies, or from abundance matching techniques that link observed galaxy properties to theoretical halo mass functions (Yang et al., 2007; Viola et al., 2015; Lim et al., 2021). However, these methods are subject to significant systematic uncertainties, particularly at low halo masses and for groups with low multiplicity, where the reliability of dynamical indicators is compromised by small number statistics and projection effects (Old et al., 2015; Robotham et al., 2011; Muldrew et al., 2012).
@@
Figure 1: Comparison of the SMF from Shark to well-established observational SMFs and SMFs produced by observational data from Van Kempen et al. (2024). The figure shows the SMF ($\log\Phi$) versus stellar mass ($\log M_{\star}$) for Shark v2.0 (blue circles) alongside canonical constraints from Baldry et al. (2012) (black dashed line) and Driver et al. (2022a) (grey dashed line). Red triangles, orange diamonds, and green crosses represent our observational data from Van Kempen et al. (2024) (SGP-2dF, G23-GAMA, and G23-2dF datasets, respectively). Error bars indicate Poisson uncertainties. Shark demonstrates excellent agreement with the established SMFs and across our observational datasets, with only minor deviations. The observational datasets exhibit a slight deficit compared to the SMFs from Baldry et al. (2012) and Driver et al. (2022a) at the low-mass end, likely attributable to WISE’s inherent limitations in detecting low-surface brightness and low dust content galaxies that contribute to the low-mass regime. The consistency between Shark v2.0 and observational constraints validates our selection of this semi-analytic model as the foundation for calibrating the baryon-halo mass relation.
@@
To provide a physically motivated and observationally calibrated framework for halo mass estimation, we chose to calibrate our halo mass estimators to Shark. This decision is motivated by Shark’s demonstrated ability to reproduce key observables, most notably the observed stellar mass function (SMF; Figure 1) and halo mass function (HMF; Figure 2), when measured using observational techniques. Furthermore, Shark is specifically tailored for application to forthcoming large-scale surveys, most notably WAVES, but also 4HS, and will serve as the primary simulation framework for training and assessing future group-finding algorithms to be used in these surveys Lagos et al. (2019). In addition, recent advancements in Shark, such as the implementation of the HBT+HERONS merger tree algorithm (Chandro-Gómez et al., 2025), have significantly reduced numerical artefacts, including: mass swapping, massive transients, and orphan galaxies—that can arise in dark matter merger trees. These improvements yield a more stable and physically consistent halo population, which is essential for the development of robust and reliable halo mass estimators.
@@
Figure 2: Comparison of the HMF derived from Shark with analytic and observational benchmarks. The blue data points represent the Shark HMF, with error bars indicating Poisson uncertainties in each mass bin. The solid black curve denotes the analytic HMF prediction for the input cosmology of Shark, computed using the hmf Python package (Murray et al., 2013). The grey dashed line shows the empirical fit from Driver et al. (2022a), based on GAMA5, SDSS5, and REFLEX II data. The magenta dashed curve corresponds to a Schechter function fit to the Shark HMF. The close agreement between the Shark HMF and the analytic prediction across the full mass range demonstrates the fidelity of Shark in reproducing the expected halo abundance for the adopted cosmology. This concordance underpins the selection of Shark as the calibration basis for our baryon–halo mass relation, ensuring that systematic biases in the halo mass distribution are minimised. Residual deviations at the low-mass end reflect both the intrinsic limitations of group identification in low-multiplicity regimes and the increasing impact of resolution effects. Overall, the robust match between Shark and the analytic HMF validates the use of Shark for calibrating observational mass proxies and for probing the baryon–dark matter connection in group-scale haloes.
@@
As shown in Figure 1, Shark provides an excellent match to well-established observational SMFs (Baldry et al., 2012; Driver et al., 2022a) across the full stellar mass range. The observational datasets from Van Kempen et al. (2024) (SGP-2dF, G23-GAMA, and G23-2dF) show good agreement with the Baldry et al. (2012) and Driver et al. (2022a) SMFs across the intermediate to high stellar mass range, which demonstrates the reliability of the sample in this regime. At lower stellar masses, however, the observed SMFs exhibit a deficit relative to these fits. This offset is attributable to the limitations of WISE photometry in detecting low-surface brightness and low stellar mass galaxies, resulting in systematic incompleteness at the faint end. Such incompleteness is a well-known limitation of near-infrared selected samples and must be considered when interpreting the low-mass behaviour of the observed SMF.
@@
where $V_{\mathrm{max}}^{i}$ is the maximum comoving volume within which the $i$th galaxy (for the SMF) or group (for the HMF) could be observed, given the survey magnitude limits and selection criteria. For the SMF, $V_{\mathrm{max}}^{i}$ is determined by the redshift range over which each galaxy remains above the survey flux limit. For the HMF, following the methodology of Driver et al. (2022a), $V_{\mathrm{max}}^{i}$ is estimated based on the $n$th brightest galaxy in each group (here, $n=3$), such that the group’s $V_{\mathrm{max}}$ is calculated based on the maximum survey volume that the $n$th brightest member remains above the survey flux limit. This approach ensures that the group sample is volume-limited with respect to its membership.
@@
3.1 Calibrating The Virial Theorem
The evolution of dark matter haloes is governed solely by gravitational forces, a regime that is well explored through detailed N-body simulations and SAMs, whereas galaxies are complex systems regulated by a multitude of baryonic processes, including gas cooling, star formation, and feedback (see review of Somerville & Davé, 2015). This fundamental distinction underpins the rationale for employing dynamical, gravity-based estimators for halo mass, which are expected to exhibit minimal dependence on the details of baryonic physics and thus provide a robust, model-independent approach to halo mass estimation.
@@ 3.1.1 The Virial Theorem
where $R$ is a radius (e.g., the maximum projected separation among group members from the median centre of the group) in units of $\mathrm{Mpc}$, and $G$ is the gravitational constant in units of $M_{\odot}^{-1}\,\mathrm{km}^{2}\,\mathrm{s}^{-2}\,\mathrm{Mpc}$, yielding a mass in $h^{-1}M_{\odot}$. This estimator is widely used in both observational and theoretical studies of galaxy groups and clusters (e.g., Carlberg et al., 1996; Eke et al., 2004; Robotham et al., 2011). While the $\frac{5}{3}$ coefficient is sometimes omitted in literature, it is physically motivated and essential for accurate mass estimates, particularly for massive halos.
The line-of-sight velocity dispersion, $\sigma_{\mathrm{los}}$, is calculated using the Gapper (gap) method (Beers et al., 1990), which is particularly robust for small group sizes and is less sensitive to outliers than standard deviation-based estimators. For a group of $N$ galaxies with ordered velocities $v_{1}<v_{2}<\ldots<v_{N}$, the Gapper estimator is defined as:
@@
where $g_{i}=v_{i+1}-v_{i}$ is the velocity gap between adjacent ordered velocities, and $w_{i}=i(N-i)$ is a weighting factor. This method provides an unbiased estimate of the velocity dispersion, especially for groups with low multiplicity, and is the standard approach in modern group catalogues (e.g., Eke et al., 2004; Robotham et al., 2011; Old et al., 2015).
In most mock haloes, the brightest galaxy is assumed to be at rest with respect to the halo centre of mass. To account for this, the velocity dispersion is further corrected by a factor of $\sqrt{N/(N-1)}$ (Eke et al., 2004; Robotham et al., 2011). The final velocity dispersion is then given by
@@
This approach ensures that the velocity dispersion estimates are unbiased and robust, even for low-multiplicity systems, and that the dominant sources of observational uncertainty are properly propagated (Robotham et al., 2011).
The top panels of Figure 3 compare the true halo masses from the simulations to those measured using observational methodology via Equation 3 on the simulated light-cones with a $z<0.1$ limit. These results demonstrate that the traditional virial theorem mass estimator is subject to substantial scatter and systematic offsets, particularly at low halo masses and for systems with low multiplicity. As shown in the lower panels, the mean offset ($\Delta$) between the estimated and true halo masses is significant across all models, reaching values as large as 2.10-2.10 dex (SAGE), 1.63-1.63 dex (GAEA), and 1.19-1.19 dex (SHARK). These biases indicate that, without calibration, the traditional estimator systematically underestimates halo masses, especially in the low-mass regime, highlighting the need for improved approaches to halo mass estimation in group environments.
Given these limitations, it is essential to develop improved approaches for halo mass estimation in group environments. In the following section, we present a calibrated virial theorem estimator designed to address these shortcomings and provide more reliable halo mass measurements.
3.1.2 MCMC Calibration
The calibration of the virial-theorem-based halo mass estimator was performed using a Bayesian framework, employing Monte Carlo Markov Chain (MCMC) sampling to explore the posterior probability distribution of the model parameters. The likelihood function was constructed to minimise the scatter in $\log M_{\mathrm{halo}}~[M_{\odot}]$ at fixed predicted halo mass, effectively minimising $\chi^{2}$ and maximising the log-likelihood. Convergence of the MCMC chains was assessed via autocorrelation analysis. Further details of the likelihood function and the Bayesian inference methodology are provided in A.
While the virial theorem provides a physically motivated starting point, its direct application to observed galaxy groups is complicated by departures from equilibrium, projection effects, and uncertainties in group membership, especially for low-multiplicity systems (Robotham et al., 2011; Muldrew et al., 2012; Old et al., 2015). To account for these effects, a scale factor $A$ is introduced to calibrate this relation, which corrects the mass estimate based on the measured velocity dispersion and projected radius of the group:
$M_{\mathrm{halo,~CVT}}=A\frac{\sigma^{2}R}{G},$ (6)
where $A$ is a function of both velocity dispersion and group radius. The calibration of $A$ is essential to ensure unbiased halo mass estimates across the full range of group multiplicities.

### v3

@@ 3.1 SHARK
Shark v2.0 (Lagos2018; Lagos2024) is an open-source, modular semi-analytic model of galaxy formation and evolution. The latest version incorporates significant advancements, including an improved treatment of angular momentum evolution, several environmental processes including ram pressure and tidal stripping, and updated feedback models. The free parameters in Shark are calibrated to reproduce the observed stellar mass function, star formation rate density, and cold gas scaling relations at $z=0$.
@@
The Shark runs analysed in this work are based on the SURFS suite of $N$-body simulations (Elahi2018), specifically medi-SURFS, which spans $(210~h^{-1}~\mathrm{Mpc})^{3}$ with $1536^{3}$ dark matter particles, yielding a particle mass of $2.21\times 10^{8}h^{-1}M_{\odot}$. Haloes, sub-haloes, and merger trees are constructed using HBT+HERONS (ChandroGomez2025). The Shark simulated light cones used in this work correspond to WAVES WIDE (North + South) light cones, totalling $\sim 1,100\,\mathrm{deg}^{2}$ in area. These light cones were constructed using the pipeline described in Lagos2019: the survey geometry and magnitude selections are built using Stingray (Chauhan2019), and the galaxy SEDs are built using ProSpect (Robotham2020). A lower stellar mass limit of $\log M_{\star}>7.5~[M_{\odot}]$ was applied to the mock light cones to ensure completeness and reliability in the resulting galaxy sample. The simulation adopts a Planck 2015 cosmology (Planck2016), with $\Omega_{\mathrm{m}}=0.3121$, $\Omega_{\Lambda}=0.6879$, $\Omega_{\mathrm{b}}=0.0491$, $h=0.6751$, $\sigma_{8}=0.8150$, and $n_{\mathrm{s}}=0.9653$. All calibrations of the halo mass relations are performed on Shark.
3.2 SAGE
@@
The Semi-Analytic Galaxy Evolution (SAGE) model (Croton2016) is a flexible, publicly available semi-analytic framework, building upon the Munich model lineage (Croton2006). SAGE incorporates detailed prescriptions for radiative cooling, star formation, stellar and AGN feedback, black hole growth, and environmental processes. The model parameters are tuned to match the observed stellar mass function and galaxy colour distributions at $z=0$.
@@
For this study, SAGE is applied to the BOLSHOI $N$-body simulation (Klypin2011), which covers a volume of $(250~h^{-1}~\mathrm{Mpc})^{3}$ with $2048^{3}$ particles, corresponding to a mass resolution of $1.35\times 10^{8}~h^{-1}~M_{\odot}$. The SAGE simulated light cones used in this work consisted of 10 lightcones, each with an area of $\sim 1,960\,\mathrm{deg}^{2}$. Haloes are identified using the ROCKSTAR phase-space halo finder (Behroozi2013a), and merger trees are constructed with the Consistent Trees algorithm (Behroozi2013b). A lower stellar mass limit of $\log M_{\star}>7.5~[M_{\odot}]$ was applied to the mock light cones to ensure completeness and reliability in the resulting galaxy sample. The adopted cosmology is based on WMAP7 (Komatsu2011), with $\Omega_{\mathrm{m}}=0.270$, $\Omega_{\Lambda}=0.730$, $\Omega_{\mathrm{b}}=0.0469$, $h=0.70$, $\sigma_{8}=0.82$, and $n_{\mathrm{s}}=0.95$. SAGE is used solely to validate the halo mass estimators without any re-tuning, thereby probing sensitivity to differing galaxy-formation prescriptions.
3.3 GAEA
@@
The Galaxy Evolution and Assembly (GAEA) semi-analytic model (DeLucia2014; Hirschmann2016; DeLucia2024) provides an independent theoretical benchmark in this work. The latest version includes updated prescriptions for AGN feedback, environmental processes affecting satellites, black hole accretion, disk instabilities, and starburst activity (Fontanot2020; DeLucia2011).
@@ 4 Halo Mass Relations
In the group and cluster regime, halo mass estimates are traditionally derived from dynamical tracers, such as the velocity dispersion of member galaxies, or from abundance matching techniques that link observed galaxy properties to theoretical halo mass functions (Yang2007; Viola2015; Lim2021). However, these methods are subject to significant systematic uncertainties, particularly at low halo masses and for groups with low multiplicity, where the reliability of dynamical indicators is compromised by small number statistics and projection effects (Old2015; Robotham2011; Muldrew2012).
@@
Figure 2: Comparison of the SMF from Shark to well-established observational SMFs and SMFs produced by observational data from VanKempen2024. The figure shows the SMF ($\log\Phi$) versus stellar mass ($\log M_{\star}$) for Shark v2.0 (blue circles) alongside canonical constraints from Baldry2012 (black dashed line) and Driver2022 (grey dashed line). Red triangles, orange diamonds, and green crosses represent our observational data from VanKempen2024 (SGP-2dF, G23-GAMA, and G23-2dF datasets, respectively). Error bars indicate Poisson uncertainties. Shark demonstrates excellent agreement with the established SMFs and across our observational datasets, with only minor deviations.
@@
To provide a physically motivated and observationally calibrated framework for halo mass estimation, we chose to calibrate our halo mass estimators to Shark. This decision is motivated by Shark’s demonstrated ability to reproduce key observables, most notably the observed stellar mass function (SMF; Figure 2) and halo mass function (HMF; Figure 3), when measured using observational techniques. Furthermore, Shark is specifically tailored for application to forthcoming large-scale surveys, most notably WAVES, but also 4HS, and will serve as the primary simulation framework for training and assessing future group-finding algorithms to be used in these surveys Lagos2019. In addition, recent advancements in Shark, such as the implementation of the HBT+HERONS merger tree algorithm (ChandroGomez2025), have significantly reduced numerical artefacts, including: mass swapping, massive transients, and orphan galaxies—that can arise in dark matter merger trees. These improvements yield a more stable and physically consistent halo population, which is essential for the development of robust and reliable halo mass estimators.
@@
Figure 3: Comparison of the HMF derived from Shark with analytic and observational benchmarks. The blue data points represent the Shark HMF, with error bars indicating Poisson uncertainties in each mass bin. The solid black curve denotes the analytic HMF prediction for the input cosmology of Shark, computed using the hmf Python package (Murray2013). The grey dashed line shows the empirical fit from Driver2022, based on GAMA5, SDSS5, and REFLEX II data. The magenta dashed curve corresponds to a Schechter function fit to the Shark HMF.
@@
As shown in Figure 2, Shark provides an excellent match to well-established observational SMFs (Baldry2012; Driver2022) across the full stellar mass range. The observational datasets from VanKempen2024 (SGP-2dF, G23-GAMA, and G23-2dF) show good agreement with the Baldry2012 and Driver2022 SMFs across the intermediate to high stellar mass range, which demonstrates the reliability of the sample in this regime. At lower stellar masses, however, the observed SMFs exhibit a deficit relative to these fits. This offset is attributable to the limitations of WISE photometry in detecting low-surface brightness and low stellar mass galaxies, resulting in systematic incompleteness at the faint end. Such incompleteness is a well-known limitation of near-infrared selected samples and must be considered when interpreting the low-mass behaviour of the observed SMF.
@@
where $V_{\mathrm{max}}^{i}$ is the maximum comoving volume within which the $i$th galaxy (for the SMF) or group (for the HMF) could be observed, given the survey magnitude limits and selection criteria. For the SMF, $V_{\mathrm{max}}^{i}$ is determined by the redshift range over which each galaxy remains above the survey flux limit. For the HMF, following the methodology of Driver2022, $V_{\mathrm{max}}^{i}$ is estimated based on the $n$th brightest galaxy in each group (here, $n=3$), such that the group’s $V_{\mathrm{max}}$ is calculated based on the maximum survey volume that the $n$th brightest member remains above the survey flux limit. This approach ensures that the group sample is volume-limited with respect to its membership.
@@
4.1 Modifying The Virial Theorem
The evolution of dark matter haloes is governed solely by gravitational forces, a regime that is well explored through detailed N-body simulations and SAMs, whereas galaxies are complex systems regulated by a multitude of baryonic processes, including gas cooling, star formation, and feedback (see review of Somerville2015b). This fundamental distinction underpins the rationale for employing dynamical, gravity-based estimators for halo mass, which are expected to exhibit minimal dependence on the details of baryonic physics and thus provide a robust, model-independent approach to halo mass estimation.
@@ 4.1.1 The Virial Theorem
where $R$ is a radius (e.g., the maximum projected separation among group members from the median centre of the group) in units of $\mathrm{Mpc}$, and $G$ is the gravitational constant in units of $M_{\odot}^{-1}\,\mathrm{km}^{2}\,\mathrm{s}^{-2}\,\mathrm{Mpc}$, yielding a mass in $h^{-1}M_{\odot}$. This estimator is widely used in both observational and theoretical studies of galaxy groups and clusters (e.g., Carlberg1996; Eke2004; Robotham2011). While the $\frac{5}{3}$ coefficient is sometimes omitted in the literature (e.g. Finn05; Evrard2008; Lau2010; Poggianti2010; Robotham2011; Alpaslan2012), it is physically motivated and essential for accurate mass estimates, particularly for massive halos.
The line-of-sight velocity dispersion, $\sigma_{\mathrm{los}}$, is calculated using the Gapper (gap) method (Beers1990), which is particularly robust for small group sizes and is less sensitive to outliers than standard deviation-based estimators. For a group of $N$ galaxies with ordered velocities $v_{1}<v_{2}<\ldots<v_{N}$, the Gapper estimator is defined as:
@@
where $g_{i}=v_{i+1}-v_{i}$ is the velocity gap between adjacent ordered velocities, and $w_{i}=i(N-i)$ is a weighting factor. Through the use of ordered velocity gaps, the gap method provides lower sampling variance than standard-deviation or bi-weight scale estimators, and is less sensitive to interlopers and has thus become a standard approach in modern group catalogues (e.g., Eke2004; Robotham2011; Old2015).
In most mock haloes, the brightest galaxy is assumed to be at rest with respect to the halo centre of mass. To account for this, the velocity dispersion is further corrected by a factor of $\sqrt{N/(N-1)}$ (Eke2004; Robotham2011). The final velocity dispersion is then given by:
@@
This approach ensures that the velocity dispersion estimates are unbiased and robust, even for low-multiplicity systems, and that the dominant sources of observational uncertainty are properly propagated (Robotham2011).
The top panels of Figure 4 compare the true halo masses from the simulations to those measured using the virial theorem via Equation 3 on the simulated light-cones with a $z<0.1$ limit. These results demonstrate that the traditional virial theorem mass estimator is subject to substantial scatter and systematic offsets, particularly at low halo masses and for systems with low multiplicity. As shown in the lower panels, the mean offset ($\Delta$) between the estimated and true halo masses is significant across all models, reaching values as large as 2.10-2.10 dex (SAGE), 1.63-1.63 dex (GAEA), and 1.19-1.19 dex (SHARK). These biases indicate that, without modification, the virial theorem systematically underestimates halo masses, especially in the low-mass regime, highlighting the need for improved halo mass estimates in group environments. Given these limitations, it is essential to develop improved approaches for halo mass estimation in group environments. In the following section, we present a calibrated virial theorem estimator designed to address these shortcomings and provide more reliable halo mass measurements.
4.1.2 Calibration of the Modified Virial Theorem
While the virial theorem provides a physically motivated starting point, its direct application to observed galaxy groups is complicated by departures from equilibrium, projection effects, and uncertainties in group membership, especially for low-multiplicity systems (Robotham2011; Muldrew2012; Old2015). To account for these effects, we introduce a modified virial theorem (MVT), which includes a scale factor, $A$, to Equation 3 and corrects the mass estimate based on the measured velocity dispersion and projected radius of the group:
$M_{\mathrm{halo,~MVT}}=A\frac{\sigma^{2}R}{G},$ (6)
where $A$ is a function of both velocity dispersion and group radius. The calibration of $A$ is essential to ensure unbiased halo mass estimates across the full range of group halo masses.
The calibration of the MVT was performed using a Bayesian framework, employing Markov Chain Monte Carlo (MCMC) sampling to explore the posterior probability distribution of the model parameters. The likelihood function was constructed to minimise the scatter in $\log M_{\mathrm{halo}}~[M_{\odot}]$ at fixed predicted halo mass, effectively minimising $\chi^{2}$ and maximising the log-likelihood. Convergence of the MCMC chains was assessed via autocorrelation analysis. Further details of the likelihood function and the Bayesian inference methodology are provided in A.